OpenStudy (rebecca1588):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. <<

pooja195 (pooja195):

Thats correct but there is another correct answer

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (rebecca1588):

C? :)

pooja195 (pooja195):

Correct good job

OpenStudy (rebecca1588):

Thanks for your help :D

HanAkoSolo (jamierox4ev3r):

@pooja195 point slope form for the win. Well done! :)